OpenStudy (anonymous):

find the domain and range of y=x^2 -2x +3

jimthompson5910 (jim_thompson5910):

Since we can plug in ANY number into x, this means that the domain is the set of real numbers. In interval notation, the domain is \[(-\infty,\infty)\] If you graph, or find the vertex, you'll see that the lowest point is at (1,2). So this means that the lowest y value is y = 2. So the range is the set of numbers that are greater or equal to 2. In interval notation, the range is \[[2,\infty)\]

OpenStudy (anonymous):

The domain is R, as x can be any real number. If you complete the square (i think it's called) you get: \[(x-1)^2+2\] From here you can see the range (y-values) are: \[[2,\ \infty)\] as \[(x-1)^2\geq 0\ for\ all\ x\]

OpenStudy (anonymous):

thx